Twelve different video games showing substance use were observed and the duration of times of game play​ (in seconds) are listed below. The design of the study justifies the assumption that the sample can be treated as a simple random sample. Use the sample data to construct a 90​% confidence interval estimate of sigma​, the standard deviation of the duration times of game play. Assume that this sample was obtained from a population with a normal distribution.

4,195

4 ,166

4 ,782

3,903

4,185   

5,039

4,678

4,726

4,714

4,881

4,712

4054
Answer #1

Here s = 393.63 and n = 12
df = 12 - 1 = 11

α = 1 - 0.9 = 0.1
The critical values for α = 0.1 and df = 11 are Χ^2(1-α/2,n-1) = 4.575 and Χ^2(α/2,n-1) = 19.675

CI = (sqrt(11*393.63^2/19.675) , sqrt(11*393.63^2/4.575))
CI = (294.33 , 610.36)